# Exponential quantum advantages in learning quantum observables from classical data

### v1

@@ Appendix B Proof of Theorem 3 and Corollary 1
Given a discrete set of polynomially describle quantum states $S=\{\ket{\psi_{\ell}}\}_{\ell}$ and measurements operators $Q=\{Q_{m}\}_{m}$. Then for every efficient (non-adaptive) algorithm $\mathcal{A}_{W}$ which learns an arbitrary unitary $W$ from query access to it on states from $S$ and measured by operators in $Q$ there exists a distribution $\mathcal{D}$ over $\{0,1\}^{n}$, a family of unitaries $\{U(\bm{x})\}_{\bm{x}}$ and a measurement $O$ for which the concept class $\mathcal{M}_{U,W,O}$ exhibits a learning advantage.
@@
  • If the first bit $x_{1}$ is 1, then on the $n_{S}$ qubit register $U(\bm{x})$ prepares the state |$\ket{\psi_{BQP}(\bm{x_{S}})}$ such that $\bra{\psi_{BQP}(\bm{x_{S}})}Z\otimes I\otimes...\otimes I\ket{\psi_{BQP}(\bm{x_{S}})}$ outputs $+1/3$ if the $n_{S}$ bitstring $\bm{x_{S}}$ belongs to an arbitrary (previously fixed) BQP complete language $\mathcal{L}$ defined over input $\bm{x_{S}}\in\{0,1\}^{n_{S}}$, while it outputs $-1/3$ otherwise. Following the same arguments used in the proof of Lemma 1, for any $\mathcal{L}\in\mathrm{BQP}$ it always exists such $U(\bm{x})$. We thus have:
@@
Regarding the unknown observable $O({\bm{\alpha}})$, we will consider a measurement operator of the kind Eq.(9). Specifically we define a controlled operator $V_{A}$, controlled by the $1+n_{Q}$ qubit register |$\ket{x_{1}x_{2}x_{3}...x_{n_{Q}+1}}$, such that $V_{A}=\sum_{0\bm{x_{Q}}}\ket{0\bm{x_{Q}}}\bra{0\bm{x_{Q}}}\otimes V(\bm{x_{Q}})$ with $\bm{x_{Q}}=x_{2}x_{3}...x_{n_{Q}+1}$ if $x_{1}=0$ and $V_{A}=\sum_{1\bm{x_{Q}}}\ket{1\bm{x_{Q}}}\bra{1\bm{x_{Q}}}\otimes I^{\otimes n_{S}}$ if $x_{1}=1$. The unitary matrices $V(\bm{x_{Q}})$ are defined such that they rotate the measurement $O$ into the observable $Q_{\bm{x_{Q}}}\in Q$ described by the bitstring $\bm{x_{Q}}=x_{2}x_{3}...x_{n_{Q}+1}$, i.e. $Q_{\bm{x_{Q}}}=V(\bm{x_{Q}})OV^{\dagger}(\bm{x_{Q}})$ where $O$ could be taken as the local observable $Z_{1}=Z\otimes I^{\otimes n_{S}-1}$ on the $n_{S}$ qubit register. The final observable $O(\bm{\alpha})$ will then be $O(\bm{\alpha})=(I^{\otimes(1+n_{Q})}\otimes W(\bm{\alpha})\;V_{A})O(V_{A}\;I^{\otimes(1+n_{Q})}\otimes W(\bm{\alpha}))^{\dagger}$, with $W({\bm{\alpha}})$ an arbitrary unitary parameterized by ${\bm{\alpha}}$. Furthermore, we define $W({\bm{\alpha}}=\vec{0})=I\otimes...\otimes I$. We can now construct a distribution $\mathcal{D}$ such that the concept class $\mathcal{M}_{U,W,O}$ exhibits a learning separation. For Lemma 1, there exists a distribution $\mathcal{D}_{\mathrm{hard}}$ for which the concept $f^{0}=\text{Tr}[\rho_{U}(\mathbf{x})V_{A}OV^{\dagger}_{A}]$ cannot be efficiently learned by a classical algorithm when $U(\bm{x})$ is the circuit which decides the BQP-complete language $\mathcal{L}$ over the $n_{S}$-sized bitstrings $\bm{x_{S}}$. We now define the following distribution $\mathcal{D}$ on the input bitsrings $\bm{x}\in\{0,1\}^{n}$:
@@
  • If $x_{1}=1$ then the following $n_{Q}$ bit $\bm{x_{Q}}$ are extracted uniformly at random while the $n_{S}$ bit $\bm{x_{S}}$ are sampled with respect to the distribution $\mathcal{D}_{\mathrm{hard}}$.
Classical hardness The classical hardness of the learning task comes directly from the proof of Lemma 1. Consider the concept $f^{0}\in\mathcal{M}_{U,W,M}$ defined by $W(\bm{0})=I^{\otimes n_{S}}$. For the same reasoning of the proof of Lemma 1 there cannot exists a polynomial sized classical circuit which evaluates $f^{0}(\bm{x})$ correctly on average with respect to the distribution $\mathcal{D}$ for the cases where $x_{1}=1$. As $x_{1}$ is equally sampled between 0 and 1, no classical algorithm can meet the learning condition of Eq. 1 on half of the input bitstrings, thus it can not learn the concept $f^{0}\in\mathcal{M}_{U,W,O}$ in polynomial time for every $\epsilon$. As the learning algorithm must suceed for every $\bm{\alpha}$, this suffices to prove the classical hardness of the learning task.
@@
We provide here the the rigorous proof of Corollary 1. The proof follows the same steps of the one for Theorem 3 above while concretizing the result for shallows $W$.
@@
Lemma 4 (Lemma 10 in [HLB+24], Learning a few-body observable with an unknown support).
@@
The proof of Lemma 1 then goes by explicitly constructing a family of parametrized circuit $\{U(\bm{x})\}_{\bm{x}}$ and an input distribution $\mathcal{D}$ such that the concept class $\mathcal{M}_{U,W,M}$ is quantum learnable using Lemma 4 while still being classically hard. Let rewrite here Corollary 1
Corollary 2 (Corollary 1 in the main text).
There exists a family of parametrized unitaries $\{U(\bm{x})\}_{\bm{x}}$ and parametrized shallow circuits $\{W({\bm{\alpha}})\}_{\bm{\alpha}}$, a measurement $O$ and a distribution $\mathcal{D}$ over $\bm{x}\in\{0,1\}^{n}$ such that the concept class $\mathcal{M}_{U,W,O}$ is not classically learnable with respect to the input distribution $\mathcal{D}$. However, there exists a quantum algorithm which learns $\mathcal{M}_{U,W,O}$ on the input distribution $\mathcal{D}$.
@@
Let consider now the case when $x_{1}=1$. Be $\mathcal{L}$ a BQP-complete language defined over the input $\tilde{\bm{x}}\in\{0,1\}^{n_{S}}$. Then we define the set of $\{U(\bm{x})\}_{\bm{x}}$ in the following way:
@@
where |$\ket{\psi_{BQP}(\bm{x})}$ is such that $\bra{\psi_{BQP}(\bm{x})}Z\otimes I\otimes...\otimes I\ket{\psi_{BQP}(\bm{x})}=1/3$ if $\tilde{\bm{x}}\in\mathcal{L}$ and $-1/3$ if $\tilde{\bm{x}}\not\in\mathcal{L}$. As we described in the Proof of Lemma 1, such a circuit $U_{BQP}(\bm{x})$ always exists for any BQP language.
As a set of unitaries $\{W(\bm{\alpha})\}_{\bm{\alpha}}$ we consider any set of parametrized unitaries acting on the $n_{S}$ qubit register such that for each $\bm{\alpha}$ parameter $W(\bm{\alpha})$ is shallow and such that $W(\bm{0})=I^{\otimes n_{S}}$. The $n_{S}$ qubit register is then measured by the observable $M=Z\otimes I\otimes...\otimes I$. Finally we define the distribution $\mathcal{D}$ on the input bitstrings $\bm{x}\in\{0,1\}^{n}$ to be the following:
@@
  • If $x_{1}=1$ then the following $n_{S}$ bits $x_{2}x_{3}...x_{n_{S}+1}$ are extracted following the distribution $\mathcal{D}_{\mathrm{hard}}$ such that $(\mathcal{L},\mathcal{D}_{\mathrm{hard}})\not\subset\mathrm{HeurP}/\mathrm{poly}$. As we assumed $\mathcal{L}\in\mathrm{BQP}$-complete such a distribution always exists by Lemma 3 in[GD23]. The other $n-(n_{S}+1)$ input bits are sampled uniformly at random.
We now show that the concept class $\mathcal{M}_{U,W,M}$ defined in Eq. 9 with $\{U(\bm{x})\}_{\bm{x}}$, $\{W(\bm{\alpha})\}_{\bm{\alpha}}$ and $M$ considered above exhibits a learning separation with respect the input distribution $\mathcal{D}$.
@@
Quantum learnability The quantum learnability is guaranteed by Lemma 4 in [HLB+24]. Recall that the training data the learning algorithm receives for a concept $f^{\bm{\alpha}}\in\mathcal{M}_{U,W,M}$ are $\mathcal{T}^{\bm{\alpha}}=\{\bm{x}_{\ell},y_{\ell}\}_{\ell}$ with $\mathbb{E}[y_{\ell}]=\text{Tr}[\rho_{U}(\bm{x}_{\ell})O({\bm{\alpha}})]$. Now, in the case the first bit of $\bm{x}_{\ell}$ is 0, $\mathcal{T}^{\bm{\alpha}}$ is exactly the training set $\mathcal{T}_{\bm{\alpha}}$ required by the algorithm in Lemma 4 to learn $O(\bm{\alpha})$. Since for every $\bm{\alpha}$ the unitary $W(\bm{\alpha})$ is of shallow depth, the locality of $O(\bm{\alpha})$ scales logarithmic with the number of qubit $n_{S}$. Then Lemma 4 guarantees that the learning algorithm runs in polynomial time requiring a polynomial-sized dataset, a condition met as half of the training samples in $\mathcal{T}^{\bm{\alpha}}$ will suffice.
{quantikz}
[…]
Figure 2: Quantum model which exhibits a learning separation for the concept class $\mathcal{M}_{U,W,O}$ when $W$ is a shallow circuit. The unitary $U(\bm{x})$ prepares the state |$\ket{\psi_{U}(\bm{x})}=\ket{x_{1}}\otimes\ket{\psi^{n_{S}}_{U}(\bm{x_{S}})}$ where the form of |$\ket{\psi^{n_{S}}_{U}(\bm{x_{S}})}$ depends on $x_{1}$, the first bit of each input $\bm{x}\in\{0,1\}^{n}$. If $x_{1}=0$, then |$\ket{\psi^{n_{S}}_{U}(\bm{x_{S}})}=\ket{\psi_{stab}(\bm{x_{S}})}=\bigotimes_{i=1}^{n_{S}}\ket{\psi^{i}_{stab1}(\bm{x_{S}})}$ is the $n_{S}=\left\lfloor{\frac{n}{3}}\right\rfloor$ qubit tensor product of single-qubit stabilizers described by the classical bitstring $x_{2}x_{3}...x_{n}$. If $x_{1}=1$, then |$\ket{\psi^{n_{S}}_{U}(\bm{x_{S}})}$ is the quantum state which decides the $n_{S}$ input bits $\bm{x_{S}}=x_{2}x_{3}...x_{n_{S}+1}$, considered as input of a BQP-complete language $\mathcal{L}$ over the bitstrings $\bm{x_{S}}\in\{0,1\}^{n_{S}}$ $W({\bm{\alpha}})$ is a parametrized shallow unitary, to prove classical hardness it is sufficient to consider $W({\bm{\alpha}}=0)=I^{\otimes n_{S}}$ and the measurement operator to be $O=Z\otimes I\otimes...\otimes I$ on the $n_{S}$ qubits register.

### v2

@@ Appendix D Proof of Theorem 5 and Corollary 1
Assuming that $\operatorname{\mathsf{BQP}}\not\subseteq\mathsf{P/poly}$, the following holds. Let $S=\{\ket{\psi_{\ell}}\}_{\ell}$ and $Q=\{Q_{m}\}_{m}$ be a discrete set of quantum states and measurement operators, respectively, both of which are polynomially describable. Then for every efficient (non-adaptive) algorithm $\mathcal{A}_{W}$ which learns an arbitrary unitary $W$ from query access to it on states from $S$ and measured by operators in $Q$ there exists a set of distributions $\{\mathcal{D}_{i}\}_{i}$ over $\{0,1\}^{n}$, a family of unitaries $\{U(\bm{x})\}_{\bm{x}}$ and a measurement $O$ for which the concept class $\mathcal{M}_{U,W,O}$ exhibits a learning advantage under the learning condition of Def. 4.
@@
  • If the first bit $x_{1}$ is 1, then on the $n_{S}$ qubit register $U(\bm{x})$ prepares the state |$\ket{\psi_{BQP}(\bm{x_{S}})}$ such that $\bra{\psi_{BQP}(\bm{x_{S}})}Z\otimes I\otimes...\otimes I\ket{\psi_{BQP}(\bm{x_{S}})}$ outputs $+1/3$ if the $n_{S}$ bitstring $\bm{x_{S}}$ belongs to an arbitrary (previously fixed) $\operatorname{\mathsf{BQP}}$complete language $\mathcal{L}$ defined over input $\bm{x_{S}}\in\{0,1\}^{n_{S}}$, while it outputs $-1/3$ otherwise. Following the same arguments used in the proof of Theorem 2, for any $\mathcal{L}\in\operatorname{\mathsf{BQP}}$ it always exists such $U(\bm{x})$. We thus have:
@@
Regarding the unknown observable $O({\bm{\alpha}})$, we will consider a measurement operator of the kind Eq.(9). Specifically we define a controlled operator $V_{A}$, controlled by the $1+n_{Q}$ qubit register |$\ket{x_{1}x_{2}x_{3}...x_{n_{Q}+1}}$, such that $V_{A}=\sum_{0\bm{x_{Q}}}\ket{0\bm{x_{Q}}}\bra{0\bm{x_{Q}}}\otimes V(\bm{x_{Q}})$ with $\bm{x_{Q}}=x_{2}x_{3}...x_{n_{Q}+1}$ if $x_{1}=0$ and $V_{A}=\sum_{1\bm{x_{Q}}}\ket{1\bm{x_{Q}}}\bra{1\bm{x_{Q}}}\otimes I^{\otimes n_{S}}$ if $x_{1}=1$. The unitary matrices $V(\bm{x_{Q}})$ are defined such that they rotate the measurement $O$ into the observable $Q_{\bm{x_{Q}}}\in Q$ described by the bitstring $\bm{x_{Q}}=x_{2}x_{3}...x_{n_{Q}+1}$, i.e. $Q_{\bm{x_{Q}}}=V(\bm{x_{Q}})OV^{\dagger}(\bm{x_{Q}})$ where $O$ could be taken as the local observable $Z_{1}=Z\otimes I^{\otimes n_{S}-1}$ on the $n_{S}$ qubit register. The final observable $O(\bm{\alpha})$ will then be $O(\bm{\alpha})=(I^{\otimes(1+n_{Q})}\otimes W(\bm{\alpha})\;V_{A})O(V_{A}\;I^{\otimes(1+n_{Q})}\otimes W(\bm{\alpha}))^{\dagger}$, with $W({\bm{\alpha}})$ an arbitrary unitary parameterized by ${\bm{\alpha}}$. Furthermore, we define $W({\bm{\alpha}}=\vec{0})=I\otimes...\otimes I$. We can now construct a set of distributions $\{\mathcal{D}_{i}\}_{i}$ such that the concept class $\mathcal{M}_{U,W,O}$ exhibits a learning separation. For Theorem 2, under the learning condition in Def. 4, the concept $f^{0}=\text{Tr}[\rho_{U}(\mathbf{x})V_{A}OV^{\dagger}_{A}]$ cannot be efficiently learned by a classical algorithm when $U(\bm{x})$ is the circuit which decides the $\operatorname{\mathsf{BQP}}$-complete language $\mathcal{L}$ over the $n_{S}$-sized bitstrings $\bm{x_{S}}$. We now define the following set distributions $\{\mathcal{D}_{i}\}_{i}$ on the input bitsrings $\bm{x}\in\{0,1\}^{n}$:
@@
  • If $x_{1}=1$ then the following $n_{Q}$ bit $\bm{x_{Q}}$ are extracted uniformly at random while the $n_{S}$ bit $\bm{x_{S}}$ are sampled with an arbitrary input distribution over $\bm{x_{S}}\in\{0,1\}^{n_{S}}$ which specifies the overall distribution $\mathcal{D}_{i}$ for each $i$.
Classical hardness The classical hardness of the learning task comes directly from the proof of Theorem 2. Consider the concept $f^{0}\in\mathcal{M}_{U,W,M}$ defined by $W(\bm{0})=I^{\otimes n_{S}}$. For the same reasoning of the proof of Theorem 2 there cannot exists a polynomial sized classical circuit which evaluates $f^{0}(\bm{x})$. Since for any of the input distributions $\mathcal{D}_{i}$, $x_{1}$ is equally sampled between 0 and 1, no classical algorithm can meet the learning condition of Eq. 4 on half of the input bitstrings, thus it can not learn the concept $f^{0}\in\mathcal{M}_{U,W,O}$ in polynomial time for every $\epsilon$. As the learning algorithm must suceed for every $\bm{\alpha}$, this suffices to prove the classical hardness of the learning task.
@@
We now provide here the the rigorous proof of Corollary 1. The proof follows the same steps of the one for Theorem 5 above while concretizing the result for shallows $W$.
@@
Lemma 3 (Lemma 10 in [34], Learning a few-body observable with an unknown support).
@@
The proof of Corollary 1 then goes by explicitly constructing a family of parametrized circuit $\{U(\bm{x})\}_{\bm{x}}$ and a set of input distributions $\{\mathcal{D}_{i}\}_{i}$ such that the concept class $\mathcal{M}_{U,W,M}$ is quantum learnable using Lemma 3 while still being classically hard. Let us rewrite here Corollary 1
See 1
@@
Let consider now the case when $x_{1}=1$. Be $\mathcal{L}$ a $\operatorname{\mathsf{BQP}}$-complete language defined over the input $\tilde{\bm{x}}\in\{0,1\}^{n_{S}}$. Then we define the set of $\{U(\bm{x})\}_{\bm{x}}$ in the following way:
@@
where |$\ket{\psi_{BQP}(\bm{x})}$ is such that $\bra{\psi_{BQP}(\bm{x})}Z\otimes I\otimes...\otimes I\ket{\psi_{BQP}(\bm{x})}=1/3$ if $\tilde{\bm{x}}\in\mathcal{L}$ and $-1/3$ if $\tilde{\bm{x}}\not\in\mathcal{L}$. As we described in the proof of Theorem 2, such a circuit $U_{BQP}(\bm{x})$ always exists for any $\operatorname{\mathsf{BQP}}$ language.
As a set of unitaries $\{W(\bm{\alpha})\}_{\bm{\alpha}}$ we consider any set of parametrized unitaries acting on the $n_{S}$ qubit register such that for each $\bm{\alpha}$ parameter $W(\bm{\alpha})$ is shallow and such that $W(\bm{0})=I^{\otimes n_{S}}$. The $n_{S}$ qubit register is then measured by the observable $M=Z\otimes I\otimes...\otimes I$. Finally we define the set of distributions $\{\mathcal{D}_{i}\}_{i}$ on the input bitstrings $\bm{x}\in\{0,1\}^{n}$ to be the following:
@@
  • If $x_{1}=1$ then the following $n_{S}$ bits $x_{2}x_{3}...x_{n_{S}+1}$ are extracted from an arbitrary distribution which characterizes $\mathcal{D}_{i}$ for each index $i$. The other $n-(n_{S}+1)$ input bits are sampled uniformly at random.
We now show that the concept class $\mathcal{M}_{U,W,M}$ defined in Eq. 9 with $\{U(\bm{x})\}_{\bm{x}}$, $\{W(\bm{\alpha})\}_{\bm{\alpha}}$ and $M$ considered above exhibits a learning separation with respect to the learning condition in Def. 4.
@@
Quantum learnability The quantum learnability is guaranteed by Lemma 3 in [34]. Recall that the training data the learning algorithm receives for a concept $f^{\bm{\alpha}}\in\mathcal{M}_{U,W,M}$ are $\mathcal{T}^{\bm{\alpha}}=\{\bm{x}_{\ell},y_{\ell}\}_{\ell}$ with $\mathbb{E}[y_{\ell}]=\text{Tr}[\rho_{U}(\bm{x}_{\ell})O({\bm{\alpha}})]$. Now, in the case the first bit of $\bm{x}_{\ell}$ is 0, $\mathcal{T}^{\bm{\alpha}}$ is exactly the training set $\mathcal{T}_{\bm{\alpha}}$ required by the algorithm in Lemma 3 to learn $O(\bm{\alpha})$. Since for every $\bm{\alpha}$ the unitary $W(\bm{\alpha})$ is of shallow depth, the locality of $O(\bm{\alpha})$ scales logarithmic with the number of qubit $n_{S}$. Then Lemma 3 guarantees that the learning algorithm runs in polynomial time requiring a polynomial-sized dataset, a condition met as half of the training samples in $\mathcal{T}^{\bm{\alpha}}$ will suffice.
Figure 2: Quantum model which exhibits a learning separation for the concept class $\mathcal{M}_{U,W,O}$ when $W$ is a shallow circuit. The unitary $U(\bm{x})$ prepares the state |$\ket{\psi_{U}(\bm{x})}=\ket{x_{1}}\otimes\ket{\psi^{n_{S}}_{U}(\bm{x_{S}})}$ where the form of |$\ket{\psi^{n_{S}}_{U}(\bm{x_{S}})}$ depends on $x_{1}$, the first bit of each input $\bm{x}\in\{0,1\}^{n}$. If $x_{1}=0$, then |$\ket{\psi^{n_{S}}_{U}(\bm{x_{S}})}=\ket{\psi_{stab}(\bm{x_{S}})}=\bigotimes_{i=1}^{n_{S}}\ket{\psi^{i}_{stab1}(\bm{x_{S}})}$ is the $n_{S}=\left\lfloor{\frac{n}{3}}\right\rfloor$ qubit tensor product of single-qubit stabilizers described by the classical bitstring $x_{2}x_{3}...x_{n}$. If $x_{1}=1$, then |$\ket{\psi^{n_{S}}_{U}(\bm{x_{S}})}$ is the quantum state which decides the $n_{S}$ input bits $\bm{x_{S}}=x_{2}x_{3}...x_{n_{S}+1}$, considered as input of a $\operatorname{\mathsf{BQP}}$-complete language $\mathcal{L}$ over the bitstrings $\bm{x_{S}}\in\{0,1\}^{n_{S}}$ $W({\bm{\alpha}})$ is a parametrized shallow unitary, to prove classical hardness it is sufficient to consider $W({\bm{\alpha}}=0)=I^{\otimes n_{S}}$ and the measurement operator to be $O=Z\otimes I\otimes...\otimes I$ on the $n_{S}$ qubits register.